Calculat the discharge if the diameter of the pipe is 30 mm with the velocity is 15 m/s
1
Expert's answer
2019-06-03T10:16:42-0400

The discharge is


"\frac{dV}{dt}=Av=0.25\pi d^2 v"

"\frac{dV}{dt}=0.25\pi (0.03)^2(15)=0.0106 \frac{m^3}{s}"